satellite73:

initial population is 500 increases by 15% in ten years

satellite73:

just like in 162 your job is to find \(A(t)=A_0e^{kt}\) \(A_0\) you are told, only need to find \(k\)

satellite73:

\[A_0=500\] the initial amount

Vrefela:

ok

satellite73:

to find \(k\) note that in 10 years you increase by 15% so have \(1.15\times 500=575\) then to find \(k\) set \[500e^{10k}=575\] and solve for \(k\)

satellite73:

note that the 500 didn't really matter because first step is to divide by 500 and get \[e^{10k}=1.15\]

satellite73:

so \[10k=\ln(1.15\] making \[k=\frac{\ln(1.15)}{10}\]

satellite73:

not that we did not mention differential equations at all

Vrefela:

oh ok

Vrefela:

so anytime there is a percent then it should be like 1.25 if 25%?

satellite73:

right

satellite73:

does this ring a bell?

Vrefela:

YES

satellite73:

now here is the diffeq part at the beginning saying the population grows at a rate proportional to the present amount is the same as saying \[\frac{dA}{dt}=kA\]

satellite73:

separating we get \[\frac{dA}{A}=kdt\] integrating we get \[\int \frac{dA}{A}=\int kt\]\[\ln(A)=kt+c\] so \[A(t)=e^{kt+c}=Ce^{kt}\]

satellite73:

we couldn't say that in 162, just took a different approach

Vrefela:

Right so its just like a separable equation?

satellite73:

exactly

satellite73:

we still didn't answer the question but now that we have the mathematical model we can do it

satellite73:

\[A(t)=500e^{.01376t}\] making \[A(30)=500e^{.01376\times 30}\]

Vrefela:

and then we solve right?

satellite73:

solve aka use a calculator

Vrefela:

Right right

satellite73:

btw if the grows rate was 8% in 12 years then \[k=\frac{\ln(1.08)}{12}\] and so on

Vrefela:

ohhh that pattern is easy

satellite73:

next

Vrefela:

What about decay? is it just a minus sign?

satellite73:

hell no

Vrefela:

oop.

satellite73:

if it decreases by 12% every 5 years then \[k=\frac{\ln(88)}{5}\]

satellite73:

actually \[\frac{\ln(.88)}{5}\]

satellite73:

because 100% - 12% = 88% = 0.88

Vrefela:

Right right. Next

satellite73:

5?

Vrefela:

Sure

satellite73:

half life is 3.3 hours start with one gram model is \[P=1\times e^{kt}\] we need k

satellite73:

in 3.3 hours you have one half of what you started with set \[e^{3.3k}=.5\] solve for \(k\)

satellite73:

so \[k=\frac{\ln(.5)}{3.3}\]

satellite73:

if the half life what 28 minutes then it would be \[k=\frac{\ln(.5)}{28}\]

satellite73:

half life means your left with half, 50% decay

satellite73:

notice again no diffeq, but the beginning part is identical to the beginning part of question 3, i see no real reason to repeat it

satellite73:

we still did not answer the question though it it clear how to proceed?

Vrefela:

So its always going to be .5?

satellite73:

for half life yes because... \(.5=\frac{1}{2}\) but only for half life

Vrefela:

lol thanks

satellite73:

do you know how to finish the problem?

Vrefela:

Plug and chug?

satellite73:

how long will it take for 90% of it to decay?

satellite73:

i am going to spend 60 seconds to get a snack when i come back you will have the answer

satellite73:

back

Vrefela:

Ce^((ln(.1))*t)

satellite73:

hell no

Vrefela:

You didn't give me parameters

satellite73:

model is \[P(t)=e^{kt}\] we found k already , what is it?

satellite73:

scroll up

Vrefela:

k=ln(.5)/28

Vrefela:

?

satellite73:

keep scrolling up

Vrefela:

k=ln(.5)/3.3

satellite73:

right what is that as a decimal?

Vrefela:

-.21

satellite73:

ok so model is \[P(t)=e^{-.21t}\] when 90%is gone what percent is left?

Vrefela:

.1

satellite73:

right solve \[e^{-.21t}=.1\] and you are done

Vrefela:

Wait so this is half life and decay?

satellite73:

same thing half life is just a specific kind of decay , where you are told how long something takes to decay by one half

Vrefela:

oh ok

satellite73:

you could decay by 10% every 8 year but half life means you decay by 50% in say 8 years

Vrefela:

oh ok

satellite73:

you still didn't answer the question

Vrefela:

oh!

satellite73:

\[e^{-.21t}=.1\] solve for \(t\)

Vrefela:

t = (ln(.1))/-.21

satellite73:

ok next

Vrefela:

One more

Vrefela:

and then i'll annoy you tomorrow

satellite73:

which one?

Vrefela:

You pick

satellite73:

9 takes two seconds it is going to be \ where \[e^{3k}=.25\] solve for k lets try another

Vrefela:

k = ln(.25)/3

Vrefela:

next

satellite73:

11 do yourself

satellite73:

look it up real easy

satellite73:

13 is newton ,but it it identical to decay with one slight difference, the quantity that decays is the difference in the temperatures, not the actual temperature

satellite73:

it is 70 degrees put in a 10 degree environment, then 1/2 a minute later it is 50 degrees translate as: the initial difference was 60 1/2 a minute later the difference is 40

Vrefela:

how is it 40?

Vrefela:

Did you solve for it?

satellite73:

you are confused let me make another thread

Vrefela:

ok